# Lifshitz critical points meet Zamolodchikov perturbation theory

### v1

@@ 1 Introduction
where $\tau$ is a privileged direction which could be an arbitrary direction in euclidean signature or time in Lorentzian signature and $i=1,\dots,D-1$. The constant $z$ is known as the dynamical critical exponent and is a key property of the model characterizing its anisotropy in the Euclidean case and its non-relativistic nature in the Lorentzian case. Clearly, for generic values of $z$, Lifshitz scalings are incompatible with rotations/boosts relating the $x_{0}$ and the $x_{i}$ directions. Such additional symmetries are only possible for special values of $z$, such as $z=1$ where we have the usual relativistic CFTs, or for $z=2$ where Galilean boosts emerge Nishida and Son (2007); Baiguera (2024); Grinstein and Pal (2018).
@@ 4.2 Coupling minimal models
It is clear from the calculation above that while the couplings and anomalous dimensions can only be trusted to order $1/m$, the leading contribution to $z$ is at order $1/m^{2}$ and this term will not get modified at higher order in the perturbative expansion. This is entirely analogous to the $c$-theorem sum rules Zamolodchikov (1986); Cardy (1988) which reliably predict an order $1/m^{3}$ modification of the central charge from an order $1/m$ calculation.
@@ 5 Conclusions and Outlook
In this note, we studied weakly relevant spin-1 perturbations that describe RG flows between CFTs and weakly anisotropic Lifshitz fixed points with $|z-1|\ll 1$. We introduced a simple system of $N=2$ coupled minimal models $\mathcal{M}_{m,m+1}\otimes\mathcal{M}_{m,m+1}$ where perturbation theory becomes reliable through the introduction of a small parameter $1/m$ and were able to compute $z$ to first non-trivial order in the associated expansion. We also observed the existence of a one-loop marginal nudge operator, [^17] In the future, it would be interesting to understand the marginality of the nudge operator more deeply and non-perturbatively, perhaps leading to sum-rules similar to those of Gabai et al. (2025); Girault et al. (2025) which rotates the frame of anisotropy and the emergence of rotational symmetry in the infrared.
It would be interesting to study the more general case of $N$ coupled minimal models, and to study these systems non-perturbatively directly on the lattice. It is straightforward to write down a Hamiltonian for statistical 2d Ising models coupled via a spin-1 operator
@@ Acknowledgements
AA is especially thankful to João Penedones for encouraging the study of Lifshitz fixed points in this context, to Gabriel Cuomo for illuminating discussions and for providing many useful references, to Adam Nahum for patient discussions on related work and to Andreia Gonçalves for continued inspiration. We further thank Apratim Kaviraj and Miguel Paulos for comments on the draft and Connor Behan, Xiangyu Cao, Subham Dutta-Chowdhury, Vasco Gonçalves, Johan Henriksson, Apratim Kaviraj, Stefanos Kousvos, Petr Kravchuk, John McGreevy, Marco Meineri, Sridip Pal, Miguel Paulos, Alessandro Podo, Giovanni Rizi, Junchen Rong, Slava Rychkov, Adar Sharon and Dam Son for useful conversations. AA acknowledges hospitality and funding from Perimeter Institute during the final stages of writing of the draft. AA is funded by the European Union (ERC, FUNBOOTS, project number 101043588). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ 1 Introduction
where $\tau$ is a privileged direction which could be an arbitrary direction in euclidean signature or time in Lorentzian signature and $i=1,\dots,D-1$. The constant $z$ is known as the dynamical critical exponent and is a key property of the model characterizing its anisotropy in the Euclidean case and its non-relativistic nature in the Lorentzian case. Clearly, for generic values of $z$, Lifshitz scalings are incompatible with rotations/boosts relating the $x_{0}$ and the $x_{i}$ directions. Such additional symmetries are only possible for special values of $z$, such as $z=1$ where we have the usual relativistic CFTs, or for $z=2$ where Galilean boosts emerge Nishida and Son (2007); Baiguera (2024); Grinstein and Pal (2018); Boisvert et al. (2025).
@@ 4.2 Coupling minimal models
It is clear from the calculation above that while the couplings and anomalous dimensions can only be trusted to order $1/m$, the leading contribution to $z$ is at order $1/m^{2}$ and this term will not get modified at higher order in the perturbative expansion. [^14] See also Korovin et al. (2013) for a closely related computation in the case of classically marginal spin-1 deformations. This is entirely analogous to the $c$-theorem sum rules Zamolodchikov (1986); Cardy (1988) which reliably predict an order $1/m^{3}$ modification of the central charge from an order $1/m$ calculation.
@@ 5 Conclusions and Outlook
In this note, we studied weakly relevant spin-1 perturbations that describe RG flows between CFTs and weakly anisotropic Lifshitz fixed points with $|z-1|\ll 1$. We introduced a simple system of $N=2$ coupled minimal models $\mathcal{M}_{m,m+1}\otimes\mathcal{M}_{m,m+1}$ where perturbation theory becomes reliable through the introduction of a small parameter $1/m$ and were able to compute $z$ to first non-trivial order in the associated expansion. We also observed the existence of a one-loop marginal nudge operator, [^18] In the future, it would be interesting to understand the marginality of the nudge operator more deeply and non-perturbatively, perhaps leading to sum-rules similar to those of Gabai et al. (2025); Girault et al. (2025). which rotates the frame of anisotropy and the emergence of rotational symmetry in the infrared.
It would be interesting to study the more general case of $N$ coupled minimal models, and to study these systems non-perturbatively directly on the lattice. It is straightforward to write down a Hamiltonian for statistical 2d Ising models coupled via a spin-1 operator [^19] In fact this model is closely related to the axial next-to-nearest-neighbor Ising model (ANNNI) Selke (1988), which was studied in the continuum via bosonization in Allen et al. (2001). We thank Philippe Lecheminant for bringing this to our attention.
@@ Acknowledgements
AA is especially thankful to João Penedones for encouraging the study of Lifshitz fixed points in this context, to Gabriel Cuomo for illuminating discussions and for providing many useful references, to Adam Nahum for patient discussions on related work and to Andreia Gonçalves for continued inspiration. We further thank Apratim Kaviraj and Miguel Paulos for comments on the draft and Connor Behan, Xiangyu Cao, Subham Dutta-Chowdhury, Vasco Gonçalves, Johan Henriksson, Apratim Kaviraj, Stefanos Kousvos, Petr Kravchuk, Philippe Lecheminant, John McGreevy, Marco Meineri, Sridip Pal, Miguel Paulos, Alessandro Podo, Giovanni Rizi, Junchen Rong, Slava Rychkov, Adar Sharon, Dam Son and Balt van Rees for useful conversations. AA acknowledges hospitality and funding from Perimeter Institute during the final stages of writing of the draft. AA is funded by the European Union (ERC, FUNBOOTS, project number 101043588). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.